Pablo Company calculates the cost for an equivalent unit of production using process costing.

Data for June
Work-in-process inventory, June 1: 25,000 units
Direct materials: 100% complete $ 50,000
Conversion: 40% complete 20,000
Balance in work-in-process, June 1 $ 70,000
Units started during June 59,000
Units completed and transferred out 59,000
Work-in-process inventory, June 30 25,000
Direct materials: 100% complete
Conversion: 80% complete
Costs incurred during June
Direct materials $ 188,800
Conversion costs
Direct labor 188,800
Applied overhead 236,000
Total conversion costs $ 424,800

Required:

1. Compute the cost per equivalent unit for both the weighted-average and FIFO methods. (Round your answers to 3 decimal places.)
Answer #1

1) Under weighted average method

Equivalent unit of material = 59000+25000 = 84000

Equivalent unit of conversion = 59000+(25000*80%) = 79000

Cost per equivalent unit of material = (50000+188800)/84000 = 2.843

Cost per equivalent unit of conversion = (20000+424800)/79000 = 5.630

2) Under FIFO

Equivalent unit of material = 59000

Equivalent unit of conversion = (25000*60%+34000+25000*80%) = 69000

Cost per equivalent unit of material = 188800/59000 = 3.20

Cost per equivalent unit of conversion = 424800/69000 = 6.157